# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ B.1.2 Larger models benefit more from AligNet
In Fig. 7 we compare supervised ImageNet-trained ViTs of different sizes. Interestingly, the smallest version of the ViTs, ViT-S, achieved the best base model performance across the board. That is, their original representation space was better aligned with the different human similarity judgments than the representation spaces of the larger ViT backbones (see Fig. 7 and Tables 12). However, after fine-tuning the models on AligNet the larger ViT versions were either better or equally-well aligned than ViT-S.
B.1.3 Detailed results for Levels
@@
Global coarse-grained. In this setting, AligNet fine-tuning improved alignment with human similarity judgments most significantly. Before fine-tuning, models achieved poor odd-one-out accuracies of $35.45\%$ (ViT-L) -- $57.38\%$ (DINOv2 ViT-B). Fine-tuning models on AligNet closed this gap and significantly improved the alignment between human and model responses to the extent that all models performed with odd-one-out accuracies of $65.70\%$ (DINOv1 ViT-B) -- $68.60\%$ (ViT-L) above the human-to-human reliability score of $61.92\%$. Interestingly, the most poorly performing model before fine-tuning, ViT-L, achieved the highest odd-one-out accuracy after fine-tuning on AligNet, with a relative improvement of almost $93.51\%$. Even the worst performing AligNet-tuned model, DINOv1 ViT-B, performed with an odd-one-out accuracy of $65.70\%$ significantly better than the best performing base model, DINOv2 ViT-B with $57.38\%$, and better than the best performing model after vanilla distillation without alignment, ViT-S with $58.82\%$, (see the coarse-grained column in Tab. 3). The relative performance improvements ranged from $19.48\%$ (DINOv2 ViT-B) -- $93.51\%$ (ViT-L). Three models—SigLIP ViT-B, DINOv2 ViT-B, and ViT-L—performed even better than the teacher model which is the strongest baseline in that setting that exists at the time of writing this manuscript.
Local fine-grained. The responses of most base models’ did not strongly correspond to human responses for fine-grained semantics either and, thus, model performances were far from the human reliabiltiy score. The human noise-ceiling was with $65.92\%$ similar to the noise-ceiling of the coarse-grained setting. Prior to fine-tuning, models achieved poor alignment scores of $40.75\%$ (ViT-L) -- $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B whose base model performance ($62.92$%) was significantly better than the performance of the other models (see Tab 3).
Soft-alignment improved this mismatch to some degree but not as significantly as it did for the coarse-grained abstraction setting. AligNet models achieved odd-one-out accuracies of $58.93$ (ViT-S) -- $62.92\%$ (DINOv1 ViT-B) (see the fine-grained column in Tab. 3). Here, relative increases in performance ranged from $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L). For a single model, DINOv1 ViT-B, the alignment score did not change after soft-alignment. Again, the best performing soft-aligned model, DINOv1 ViT-B, performed with an odd-one-out accuracy of $62.92\%$ better than any other model—whether linearly aligned using UD or fine-tuned via teacher distillation—in that setting. Note that UD and fine-tuning via distillation decreased the performance of DINOv1 ViT-B. This model performed closest to the human reliability score of $65.92\%$. Every soft-aligned model performed better than the teacher in that abstraction setting.
@@
AligNet fine-tuning changed the representation spaces of all models to be equally well aligned with the human respondents. The differences between AligNet models was significantly smaller compared to the other settings, with odd-one-out accuracies of $93.09\%-94.24\%$. Surprisingly, the odd-one-out accuracies of models fine-tuned on AligNet were higher than the human reliability score of $89.21\%$ (see Fig. 2D, rightmost column) with the best model achieving an impressive alignment score of $94.24\%$ (ViT-L). This means that the responses of AligNet models were more similar to the average human responses—since each triplet response is the majority response of the subject population—than the level of agreement among the human subjects themselves. The performances of all models improved by fine-tuning them on AligNet. The relative increases in performance were between $0.62\%-32.39\%$.
@@
B.1.4 Correlation of model output uncertainties with human RTs
@@
Here, we evaluate how the uncertainties of the model outputs—measured as the discrete Shannon entropy of the triplet probabilities—correlate with the human response times for the different triplets. For all base models, we observed a poor (close to zero) Spearman rank correlation between its output uncertainties over the triplet odd-one-out choices and the human response times for both the global coarse-grained and the local fine-grained semantic settings (see panel A in Fig. 8). However, except for ViT-B and ViT-L, most base models showed a medium positive Spearman rank correlation ($\rho$=0.3-0.4) for the class-boundary setting (see rightmost column in panel A Fig. 8). Note that the class-boundary setting is the easiest setting with the least variance/disagreement among the human participants (see Sec. A.3 for details). AligNet-finetuning significantly improved the correspondence between model output uncertainties and human responses times for all models and triplet type settings. The improvements were most striking for the global coarse-grained semantic setting where every model achieved a medium to strong positive Spearman rank correlation of close to $\rho$=0.4. For the class-boundary setting, models fine-tuned on AligNet even achieved a strong positive Spearman rank correlation coefficient of close to $\rho=0.5$.
@@
In panel B of Fig. 8, we compare AligNet fine-tuned representations (darker colors) against performing distillation without alignment (lighter colors). We can see that simply performing distillation without alignment could improve the correspondence of model uncertainties and human response times over the base representations but the improvement was not as substantial compared to applying the full AligNet framework. We find that for every model and every triplet setting, AligNet fine-tuning yielded a significantly stronger Spearman rank correlation coefficient. In Fig. 9, Fig. 10, and Fig. 11 we contrast the human response times (in $\log$-space) against the model’s output uncertainties over the triplet odd-one-out responses for ViT-B, DINOv2 ViT-B, and SigLIP ViT-B respectively for each triplet individually. We chose those three models because they reflect a representative subset of the student models (see Sec. A.1.5 that we fine-tuned on AligNet. Each of those models has the same backbone but was pretrained on a different datasets with a different objective function. We observe that AligNet fine-tuning significantly improved the correspondence of the models’ output uncertainties to the human responses times for all three models and triplet settings. Interestingly, distillation without alignment, decreased the Spearman rank correlation coefficient of DINOv2 for each of the three triplet settings (see panel B in Fig. 10). That was not the case for ViT-B and SigLIP ViT-B (see panel B in Fig. 9 and Fig. 11 respectively).
@@ B.2.1 Few-shot learning
Few-shot learning [Wang et al., 2020] is a common way to evaluate how well neural network representations generalize to new tasks: Here, the goal is to learn a new task/dataset from only a few labelled examples (typically $\leq$10 examples per labelled class are used to train a classifier on top of the pretrained representations), and then evaluate its performance on the full test set for that task. Good few-shot performance is indicative of models that have a solid understanding of many different concepts, and that their representations are able to quickly integrate/expand to new or related concepts.
We run few-shot learning evaluations on eight different supervised learning tasks. ImageNet [Russakovsky et al., 2015], Places365 [Zhou et al., 2018] and Cifar100 [Krizhevsky and Hinton, 2009] test for broad understanding of a wide variety of visual concepts, and gauge how easily a representation can adapt to a new learning task on concepts a (pretrained) model has likely encountered before. We also test datasets that instead require that the model is able to differentiate between very fine-grained differences in related concepts. Concretely, we test datasets such as Caltech-UCSD Birds 200 (CUB-200) Birds [Wah et al., 2011], Stanford Cars [Krause et al., 2013], Flowers [Nilsback and Zisserman, 2008], Oxford-IIIT Pets [Parkhi et al., 2012] and the UC Merced Land Use Dataset [Yang and Newsam, 2010]. For our evaluation, we freeze the model weights, and train a linear classifier on top of the extracted image representation [Alain and Bengio, 2017]. On all datasets, we test two few-shot settings: the extreme case where only a single example per class is provided (‘1-shot‘) and the more conservative, but still challenging ‘10-shot‘ challenge.
In Fig. 12, we compare the few-shot performance of pre-trained models vs. their few-shot performance after being aligned-finetuned. AligNet finetuning is clearly beneficial on most datasets in both 1-shot and 10-shot evaluations. Even models that have previously been trained on ImageNet data (e.g. DINOv1 ViT-B/16 and supervised ViT-B/16) show improvements on 1-shot ImageNet performance, which indicates that the benefits do not merely come from being exposed to ImageNet data, but also from the label information we distilled into AligNet.
@@
Muttenthaler et al. [2023b] have demonstrated that linearly transforming the pretrained representations of state-of-the-art image/text models into a human-aligned concept space can improve upon methods that were specifically designed for improving the performance of pretrained representations in few-shot learning tasks. One of these methods is Tip-Adapter [Zhang et al., 2022]. Tip-Adapter is designed to yield effective few-shot classifiers for image/text models by linearly combining the output of two modules — a zero-shot classifier and a key-value cache model. UD (see Appx. A.1.1) and gLocal [Muttenthaler et al., 2023b] have highly similar learning objectives: both linearly transform the representations of pretrained vision foundation models into a (global) human-aligned concept space while preserving the models’ nearest neighbor structure. Since we find that UD either performs equally well or better in matching the human similarity judgments and predicting the human RTs for the triplets in the Levels data (see Appx. B.2.3) and soft-aligned models outperform the UD transformation across the different human object similarity tasks that we evaluated (see Appx. B.1 for details), we can be confident that soft-aligned image/text models (i.e., image/text models fine-tuned on AligNet) can improve upon Tip-Adapter. For CIFAR-100, the improvements in zero-shot performance of using gLocal in conjunction with Tip-Adapter compated to Tip-Adapter alone ranged from $6.29\%$ for a CLIP-RN50 up to $18.72\%$ for a CLIP ViT-L/14 (see Appendix D in Muttenthaler et al. [2023b]). For other datasets, the improvements in zero-shot performance were similarly high. One caveat with Tip-Adapter is that its application is restricted to models that have a text encoder module, i.e., image/text models. Soft-alignment, on the other hand, does not suffer from the same constraint and can be applied to any vision model.

### v4

@@ B.1.2 Larger models benefit more from AligNet
In Fig.˜7 we compare supervised ImageNet-trained ViTs of different sizes. Interestingly, the smallest version of the ViTs, ViT-S, achieved the best base model performance across the board. That is, their original representation space was better aligned with the different human similarity judgments than the representation spaces of the larger ViT backbones (see Fig.˜7 and Tables 12). However, after fine-tuning the models on AligNet the larger ViT versions were either better or equally-well aligned than ViT-S.
B.1.3 Descriptive statistics for human response times and uncertainty measures on Levels
We examined the human response time (RT) differences across the three abstraction settings as an indicator of cognitive processing demands during odd-one-out selection (see Fig.˜5A). As expected, responses were fastest for the class-boundary triplets (Mdn = 2.04 s, SE = 0.03), significantly slower for fine-grained semantic triplets (Mdn = 3.07 s, SE = 0.05; $t(447)=4{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.920}$, $p<0.001$), and slowest for coarse-grained semantic triplets (Mdn = 3.25 s, SE = 0.06; $t(447)=1{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.160}$, $p<0.001$ compared to fine-grained). Unlike the simpler choices enabled by clear semantic boundaries, longer RTs at higher levels of semantic abstraction likely reflected more complex decisions, where participants may have weighed information across multiple perceptual and semantic dimensions. Consistent with these RT results, we also found higher levels of cross-participant agreement (see Fig.˜5C) in the class boundary condition compared to both the coarse-grained semantic and the fine-grained semantic conditions.
Participants’ RTs were used as a proxy for decision uncertainty. RTs also correlated positively with cross-participant disagreement within each abstraction levels (coarse-grained semantic: $r=0.325$, $p<0.001$; fine-grained semantic: $r=0.364$, $p<0.001$; class-boundary: $r=0.443$, $p<0.001$; see SI Sec. A.3 and Fig.˜5B), suggesting that human RTs provided a reasonable approximation of decision uncertainty for model-to-human comparison.
B.1.4 Detailed results for Levels
@@
Global coarse-grained. In this setting, AligNet fine-tuning improved alignment with human similarity judgments most significantly. Before fine-tuning, models achieved poor odd-one-out accuracies of $35.45\%$ (ViT-L) - $57.38\%$ (DINOv2 ViT-B). Fine-tuning models on AligNet closed this gap and significantly improved the alignment between human and model responses to the extent that all models performed with odd-one-out accuracies of $65.70\%$ (DINOv1 ViT-B) - $68.60\%$ (ViT-L) above the human-to-human reliability score of $61.92\%$. Interestingly, the most poorly performing model before fine-tuning, ViT-L, achieved the highest odd-one-out accuracy after fine-tuning on AligNet, with a relative improvement of almost $93.51\%$. Even the worst performing AligNet-tuned model, DINOv1 ViT-B, performed with an odd-one-out accuracy of $65.70\%$ significantly better than the best performing base model, DINOv2 ViT-B with $57.38\%$, and better than the best performing model after vanilla distillation without alignment, ViT-S with $58.82\%$, (see the coarse-grained column in Tab. 3). The relative performance improvements ranged from $19.48\%$ (DINOv2 ViT-B) - $93.51\%$ (ViT-L). Three models—SigLIP ViT-B, DINOv2 ViT-B, and ViT-L—performed even better than the teacher model which is the strongest baseline in that setting that exists at the time of writing this manuscript.
Local fine-grained. The responses of most base models’ did not strongly correspond to human responses for fine-grained semantics either and, thus, model performances were far from the human reliabiltiy score. The human noise-ceiling was with $65.92\%$ similar to the noise-ceiling of the coarse-grained setting. Prior to fine-tuning, models achieved poor alignment scores of $40.75\%$ (ViT-L) - $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B whose base model performance (62.9262.92%) was significantly better than the performance of the other models (see Tab 3).
Soft-alignment improved this mismatch to some degree but not as significantly as it did for the coarse-grained abstraction setting. AligNet models achieved odd-one-out accuracies of $58.93$ (ViT-S) - $62.92\%$ (DINOv1 ViT-B) (see the fine-grained column in Tab. 3). Here, relative increases in performance ranged from $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L). For a single model, DINOv1 ViT-B, the alignment score did not change after soft-alignment. Again, the best performing soft-aligned model, DINOv1 ViT-B, performed with an odd-one-out accuracy of $62.92\%$ better than any other model—whether linearly aligned using UD or fine-tuned via teacher distillation—in that setting. Note that UD and fine-tuning via distillation decreased the performance of DINOv1 ViT-B. This model performed closest to the human reliability score of $65.92\%$. Every soft-aligned model performed better than the teacher in that abstraction setting.
@@
AligNet fine-tuning changed the representation spaces of all models to be equally well aligned with the human respondents. The differences between AligNet models was significantly smaller compared to the other settings, with odd-one-out accuracies of $93.09\%-94.24\%$. Surprisingly, the odd-one-out accuracies of models fine-tuned on AligNet were higher than the human reliability score of $89.21\%$ (see Fig.˜2D, rightmost column) with the best model achieving an impressive alignment score of $94.24\%$ (ViT-L). This means that the responses of AligNet models were more similar to the average human responses—since each triplet response is the majority response of the subject population—than the level of agreement among the human subjects themselves. The performances of all models improved by fine-tuning them on AligNet. The relative increases in performance were between $0.62\%-32.39\%$.
@@
B.1.5 Human alignment depends on the abstraction level
We found that the best model before soft-alignment at a particular abstraction level was often not the best model after soft-alignment. While DINOv2 was the best aligned model before AligNet fine-tuning for the fine-grained (57.38%) and coarse-grained (57.72%) abstraction settings (see Tab. 3), it remained the best aligned model only for the fine-grained setting (62.24%) but did not improve as significantly as the other models for the coarse-grained (relative improvement of 18.78%) and class-boundary settings (relative improvement of 4.13%). On the other hand, ViT-L was the worst-aligned model before fine-tuning for the coarse-grained setting (35.45%), but after AligNet fine-tuning, it became the best-aligned model, achieving the highest odd-one-out accuracy (68.60%) across all models. The relative improvement in performance was 93.51%. In addition, ViT-L was both the best aligned model before (93.67%) and after (94.24%) AligNet fine-tuning for the class-boundary setting (both scores are higher than the human reliability score of 89.21%).
We observed a similar phenomenon for the coarse- and fine-grained human responses datasets of Peterson et al. [62, 63]. DINOv2 showed the highest Spearman rank correlation coefficient with the human similarity judgments prior to any fine-tuning for both abstraction levels but remained the best aligned model only for the fine-grained datasets after fine-tuning (see Tab. 2 for details). This suggests that DINOv2 fine-tuned on AligNet best captured fine-grained human semantic structure whereas image/text contrastive and supervised models best reflected coarse-grained and class-boundary semantic structure respectively after the fine-tuning process.
B.1.6 ImageNet vs. Ecoset models on Levels
Here, we evaluate vision models trained on Ecoset [51] on the Levels data. Ecoset is an ecologically motivated natural image dataset, designed in the hope to produce models that better reflect human perception than models trained on ImageNet. The number of (training) data points is the same between Ecoset and ImageNet (approx. 1.3M) but the number of classes is differs between the two datasets: 565 (Ecoset) vs. 1000 (ImageNet). We find that models trained on Ecoset are worse aligned with human similiarty judgments than the same set of models trained on ImageNet in most abstraction settings (see Tab. 4). This is in line with previous findings of Muttenthaler et al. [55] who found that Ecoset models are worse aligned than ImageNet models with the human similarity judgment datasets from Hebart et al. [28], King et al. [37], and Cichy et al. [14]. The reason for this is most likely the higher number of classes in ImageNet which has been found to play a crucial role for the alignment of neural network representations with human semantic cognition [55]. Note that Ecoset is designed for supervised learning, and weakly-/self-supervised image/text model representations generally correspond better to human behavior and brain data than the representations of supervised ImageNet models [40, 55].
Coarse-grained Fine-grained Class-boundary
Model $\setminus$ Data Ecoset ImageNet Ecoset ImageNet Ecoset ImageNet
AlexNet 42.83% 38.41% 44.39% 44.31% 55.47% 49.81%
VGG-16 42.65% 40.18% 45.16% 49.95% 60.34% 64.58%
ResNet-50 44.98% 46.46% 44.80% 49.98% 63.85% 91.90%
Inception v3 45.86% 36.08% 49.59% 35.76% 70.38% 82.28%
Table 4: Human alignment results for Ecoset vs. ImageNet models on the Levels data. Here, we show triplet odd-one-out accuracies (in %)—measured as the fraction of triplets for which models selected the same odd-one-out image as the majority of the human participants for the three levels of abstraction—coarse-grained, fine-grained, and class-boundary—in the Levels dataset that we collected via online crowdsourcing. Representations of Ecoset models were extracted with the Python toolbox thingsvision [53].
B.1.7 Evaluating vision-language models on Levels
Although we focus on vision models in the main text, a growing area of research focuses on Vision-Language Models (VLMs). As language captures much of the detail of human semantic knowledge, it is natural to ask how these models perform on the Levels dataset. We therefore evaluated Gemini 2.0 Flash [77] and Gemini 2.5 Pro [24] on Levels, using a prompt that approximately follows the human instructions, except edited to describe a single trial rather than a full experimental procedure: “I will show you three object images presented in sequence. These images show different things. Your task is to pick one image that is least similar to the other two. If you do not recognize what is shown to you in one of the images, simply base your judgement on your best guess at what the image might show. Which of these images is least similar to the others? Please give your answer ("first", "second", or "third") before explaining.”
After this prompt, we presented the three images, and then evaluated the model’s response. As expected, the models—and especially the stronger 2.5 Pro model—perform fairly well at the class-boundary judgments, and moderately well at the others; they are comparable to or better than the stronger vision models before alignment. However, neither VLM achieves the performance of AligNet, and they are especially weak on the coarse-grained judgments. These results suggest that the improvements offered by our methods are complementary to the benefits offered by the richer language supervision in VLMs.
Model $\setminus$ Triplet type Coarse-grained Fine-grained Class-boundary
Gemini 2.0 Flash 55.0% 58.8% 89.2%
Gemini 2.5 Pro 61.3% 61.3% 89.9%
Table 5: Gemini 2.0 Flash and Gemini 2.5 Pro evaluation performance (odd-one-out accuracy) on the Levels dataset. The models perform well on most categories, comparable to the original versions of some of the stronger vision models—with the stronger model even outperforming them on coarse-grained triplets—but neither VLM achieves as high of performance as AligNet, especially on the coarse-grained triplets. (cf. Tab.˜3.)
B.1.8 Correlation of model output uncertainties with human RTs
@@
Here, we evaluate how the uncertainties of the model outputs—measured as the discrete Shannon entropy of the triplet probabilities—correlate with the human response times for the different triplets. For all base models, we observed a poor (close to zero) Spearman rank correlation between its output uncertainties over the triplet odd-one-out choices and the human response times for both the global coarse-grained and the local fine-grained semantic settings (see panel A in Fig.˜8). However, except for ViT-B and ViT-L, most base models showed a medium positive Spearman rank correlation ($\rho$=0.3-0.4) for the class-boundary setting (see rightmost column in panel A Fig.˜8). Note that the class-boundary setting is the easiest setting with the least variance/disagreement among the human participants (see Sec. A.3 for details). AligNet-finetuning significantly improved the correspondence between model output uncertainties and human responses times for all models and triplet type settings. The improvements were most striking for the global coarse-grained semantic setting where every model achieved a medium to strong positive Spearman rank correlation of close to $\rho$=0.4. For the class-boundary setting, models fine-tuned on AligNet even achieved a strong positive Spearman rank correlation coefficient of close to $\rho=0.5$.
@@
In panel B of Fig.˜8, we compare AligNet fine-tuned representations (darker colors) against performing distillation without alignment (lighter colors). We can see that simply performing distillation without alignment could improve the correspondence of model uncertainties and human response times over the base representations but the improvement was not as substantial compared to applying the full AligNet framework. We find that for every model and every triplet setting, AligNet fine-tuning yielded a significantly stronger Spearman rank correlation coefficient. In Fig.˜9, Fig.˜10, and Fig.˜11 we contrast the human response times (in $\log$-space) against the model’s output uncertainties over the triplet odd-one-out responses for ViT-B, DINOv2 ViT-B, and SigLIP ViT-B respectively for each triplet individually. We chose those three models because they reflect a representative subset of the student models (see Sec. A.1.5 that we fine-tuned on AligNet. Each of those models has the same backbone but was pretrained on a different datasets with a different objective function. We observe that AligNet fine-tuning significantly improved the correspondence of the models’ output uncertainties to the human responses times for all three models and triplet settings. Interestingly, distillation without alignment, decreased the Spearman rank correlation coefficient of DINOv2 for each of the three triplet settings (see panel B in Fig.˜10). That was not the case for ViT-B and SigLIP ViT-B (see panel B in Fig.˜9 and Fig.˜11 respectively).
@@ B.2.1 Few-shot learning
Few-shot learning [83] is a common way to evaluate how well neural network representations generalize to new tasks: Here, the goal is to learn a new task/dataset from only a few labelled examples (typically $\leq$10 examples per labelled class are used to train a classifier on top of the pretrained representations), and then evaluate its performance on the full test set for that task. Good few-shot performance is indicative of models that have a solid understanding of many different concepts, and that their representations are able to quickly integrate/expand to new or related concepts.
We run few-shot learning evaluations on eight different supervised learning tasks. ImageNet [73], Places365 [89] and Cifar100 [43] test for broad understanding of a wide variety of visual concepts, and gauge how easily a representation can adapt to a new learning task on concepts a (pretrained) model has likely encountered before. We also test datasets that instead require that the model is able to differentiate between very fine-grained differences in related concepts. Concretely, we test datasets such as Caltech-UCSD Birds 200 (CUB-200) Birds [82], Stanford Cars [41], Flowers [58], Oxford-IIIT Pets [61] and the UC Merced Land Use Dataset [84]. For our evaluation, we freeze the model weights, and train a linear classifier on top of the extracted image representation [2]. On all datasets, we test two few-shot settings: the extreme case where only a single example per class is provided (‘1-shot‘) and the more conservative, but still challenging ‘10-shot‘ challenge.
In Fig.˜12, we compare the few-shot performance of pre-trained models vs. their few-shot performance after being aligned-finetuned. AligNet finetuning is clearly beneficial on most datasets in both 1-shot and 10-shot evaluations. Even models that have previously been trained on ImageNet data (e.g. DINOv1 ViT-B/16 and supervised ViT-B/16) show improvements on 1-shot ImageNet performance, which indicates that the benefits do not merely come from being exposed to ImageNet data, but also from the label information we distilled into AligNet.
@@
Muttenthaler et al. [56] have demonstrated that linearly transforming the pretrained representations of state-of-the-art image/text models into a human-aligned concept space can improve upon methods that were specifically designed for improving the performance of pretrained representations in few-shot learning tasks. One of these methods is Tip-Adapter [88]. Tip-Adapter is designed to yield effective few-shot classifiers for image/text models by linearly combining the output of two modules — a zero-shot classifier and a key-value cache model. UD (see SI. A.1.1) and gLocal [56] have highly similar learning objectives: both linearly transform the representations of pretrained vision foundation models into a (global) human-aligned concept space while preserving the models’ nearest neighbor structure. Since we find that UD either performs equally well or better in matching the human similarity judgments and predicting the human RTs for the triplets in the Levels data (see SI. B.2.7) and soft-aligned models outperform the UD transformation across the different human object similarity tasks that we evaluated (see SI. B.1 for details), we can be confident that soft-aligned image/text models (i.e., image/text models fine-tuned on AligNet) can improve upon Tip-Adapter. For CIFAR-100, the improvements in zero-shot performance of using gLocal in conjunction with Tip-Adapter compated to Tip-Adapter alone ranged from $6.29\%$ for a CLIP-RN50 up to $18.72\%$ for a CLIP ViT-L/14 (see Appendix D in Muttenthaler et al. [56]). For other datasets, the improvements in zero-shot performance were similarly high. One caveat with Tip-Adapter is that its application is restricted to models that have a text encoder module, i.e., image/text models. Soft-alignment, on the other hand, does not suffer from the same constraint and can be applied to any vision model.